OpenStudy (mayankdevnani):

A parachutist after bailing out, falls 50m without friction.When parachute opens , it decelerates at 2m/s^2. He reaches the ground with a speed of 3m/s.At what height did he bail out?

OpenStudy (mayankdevnani):

options are- a) 293m b) 111m c)91m d)182m

OpenStudy (mayankdevnani):

any body........plz help me

mathslover (mathslover):

NSTSE?

OpenStudy (mayankdevnani):

oh!! exactly....

OpenStudy (anonymous):

Lets say he jumped out at height h

OpenStudy (mayankdevnani):

ok

mathslover (mathslover):

\[\large{v^2=u^2+2as}\]

OpenStudy (anonymous):

2 x 10 x 50 = v^2 v=10root10 m/s this is the downward speed when the parachute opens

mathslover (mathslover):

^ use the above equation.

OpenStudy (mayankdevnani):

ok

OpenStudy (anonymous):

now he has to fall (h-50) metres...use the same equation again

mathslover (mathslover):

since u = 0 m/s \[\large{v^2=2as}\] s = 50 m g = 9.8 m/s^2

OpenStudy (mayankdevnani):

ok then

OpenStudy (anonymous):

only this time a = 10 - 2 m/s^2 because the parachute creates a drag force in the upward direction

OpenStudy (mayankdevnani):

then

OpenStudy (mayankdevnani):

but a is already given @him1618

mathslover (mathslover):

\[\large{v=\sqrt{20 * 50} = \sqrt{1000} =10\sqrt{10} ms^{-1}}\]

OpenStudy (anonymous):

but the net acceleration is acceleration duwe to g - accn due to parachute

OpenStudy (anonymous):

or maybe u can take a as 2 m/s^2 try both..it only depends on the question..u should know the method

OpenStudy (mayankdevnani):

ok

OpenStudy (mayankdevnani):

so @him1618

mathslover (mathslover):

\[\large{10\sqrt{10}-3=u}\] \[\large{v=0}\]

OpenStudy (mayankdevnani):

@Harkirat

mathslover (mathslover):

\[\large{0=(10\sqrt{10}-3)^2+2(2)(s)}\]

mathslover (mathslover):

\[\large{0=818.6-4s}\] \[\large{4s=818.6}\] \[\large{s=24.65 m}\]

mathslover (mathslover):

sorry it is s = 204.65 m

mathslover (mathslover):

\[\large{s=204.65 m}\]

mathslover (mathslover):

add 50 m now ... 50 + 204.65 = 254.65 m

OpenStudy (mayankdevnani):

so in options it is not given

OpenStudy (mayankdevnani):

hey!! mathslover can we meet pathak sir tomorrow...

mathslover (mathslover):

Nahin ruk.

OpenStudy (mayankdevnani):

what about @mathslover done...

mathslover (mathslover):

is a) the answer?

OpenStudy (anonymous):

If you want to learn then dont go asking on websites...its not a good way to learn

mathslover (mathslover):

\[\large{0=(10\sqrt{10})^2-4s}\] \[\large{0=1000-4s}\] \[\large{4s=1000}\] \[\large{s=250 m}\] s + 50 m = 250 m + 50 m = 300 m

mathslover (mathslover):

@mayankdevnani confirm me that is a) the answer?

OpenStudy (mayankdevnani):

yaa

OpenStudy (mayankdevnani):

@mathslover

OpenStudy (mayankdevnani):

plzz

OpenStudy (experimentx):

didn't you guys get this?

OpenStudy (ajprincess):

\(v^2=u^2+2as\) \(=2*9.8*50\) \(=980\) This is the velocity of the man when the parachute opens. using \(v^2=u^2+2as\) again. \(v^2=u^2+2as\) \(9=980-2*2*s\) \(4s=980-9\) \(4s=971\) \(s=971/4\) \(s=242.75\) \(s=243\) So total height h=243+50 =293. am I right @experimentX?

OpenStudy (experimentx):

huh ... let ma have a look!!

OpenStudy (ajprincess):

oh k.

OpenStudy (experimentx):

Hmm ... well the height should be less than 50. @ajprincess give another try.

OpenStudy (anonymous):

she's right.

OpenStudy (experimentx):